OpenStudy (anonymous):

How many solutions will this system of equations have? y = 3.5x - 3.5 y = -3.5x + 3.5 No solution Infinite solutions One solution Two solutions

OpenStudy (alexandervonhumboldt2):

if you grath it only one solution

OpenStudy (anonymous):

Thank you can you help me on another one?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system? 6x + 12y = 13 2x + y = -6 x - 2y = 7 8x - 4y = -5

OpenStudy (alexandervonhumboldt2):

If the two lines are parallel, the system is inconsistent, with no solution. If the two lines intersect at one point, the system is consistent and independent, with one solution. If the two lines are the same, the system is consistent and dependent, with infinitely many solutions.

OpenStudy (alexandervonhumboldt2):

incostienent= C

OpenStudy (anonymous):

Ok thank you so much! I gave you a medal :)

OpenStudy (alexandervonhumboldt2):

u ii am not very sure about the last one